OpenStudy (anonymous):

Give a formula for a function whose domain is all negative values of x except x=-5

OpenStudy (anonymous):

Somthing like this ? f(x)= 1/(x+5) + 1/(1 - x/|x|)

OpenStudy (anonymous):

Does it make sense ?

OpenStudy (anonymous):

how did you come up with it?

OpenStudy (anonymous):

well it has two parts. This part is undefined when x=-5 (you get a division by 0): 1/(x+5) This part is undefined for non-negative numbers: 1/(1 - x/|x|) notice that |x| means absolute value of x, so: x/|x| will be 1 for postive values of x, -1 for negative values of x and undefined for x=0 only for negative numbers we wont have a division by zero below: 1/(1 - x/|x|) Let me know if this is clear :)

OpenStudy (anonymous):

this is clear. but there are no simpler equations to use?

OpenStudy (anonymous):

f(x)= 1/(x+5) + log( -x )

OpenStudy (anonymous):

f(x) = log(-x) / (x+5)